A hydrogen atom is in the 2p state. Energy is -3.40 eV

A) determine its angular momentum

B) determine its quantum number

C) determine the possible possible values of its magnetic quantum number
Answer #1

a) angular momentum = nh/2\pi   where h is planks constant and n=2 for the second exciated state

= 2*6.63*10^-34/2*\pi

= 2.1114*10^-34 kg/sec.

b) quantum number of hydrogen atom in 2p state

n l m

2 1 -1

2 1 0

2 1 1

c) When a magnetic field is applied, the 2 p level of atomic hydrogen is split into three different energy states with energy difference of \DeltaE = \muB*B* \Deltam ℓ where m ℓ = magnetis quantum number.

m ℓ = 1, 0, -1 are magnetic quantum numbers.